OpenStudy (anonymous):

(t+5)(t-5)/9t+3)(t+5)/(t-5)/(t+9)

OpenStudy (anonymous):

\[\frac{(t+5)(t-5)}{(t+3)(t+5)}\times \frac{t+9}{t-5}\] is a start

OpenStudy (anonymous):

i am assuming the 9 in the first part of the denominator is a typo and you meant \[(t+3)\] not \[9t+3)\]

OpenStudy (anonymous):

then cancel common factors and you will have your answer